# by Laurent Busé, Thomas Dedieu

### v2

@@
Abstract
The goal of this text is to understand and prove a formula stated by Salmon, which gives the first terms of some Taylor expansion of the discriminant of a plane algebraic curve. Salmon uses his formula to derive various enumerative quantities for surfaces in $\mathbf{P}^{3}$. We provide complete proofs of this formula and its enumerative applications, and extend Salmon’s considerations to hypersurfaces in a projective space of arbitrary dimension. To this end, we introduce the concept of reduced discriminant, and provide a thorough study of its weight properties; the latter are deeply linked to projective enumerative geometric properties.
Note. This text is one of a series issued from the seminar Degenerations and enumerations of curves on surfaces, held in the University of Roma Tor Vergata 2015–2017.111Seminar held in the framework of project FOSICAV, which has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 652782. We have kept for the most part the text format of this series of notes, which at times may appear odd on one text taken individually.
@@ Abstract
[…]
In his wonderful book [18], first published in 1862, Salmon casually gives the leading term of some Taylor expansion of the discriminant of a plane algebraic curve. Consider the curve $\mathrm{V}(F)\subseteq\mathbf{P}^{2}$; in a suitable system of homogeneous coordinates, its equation $F$ has the form
where each polynomial $F_{k}(X,Y)$ is homogeneous of degree $k$ in $X,Y$. Salmon states, without any explanation or reference, that the discriminant of $F$ is of the form
[…]
@@
where $\Phi$ is “the discriminant when $T$ vanishes” [18, § 605]. This note arose as an attempt to understand and prove this formula.
[…]
@@
The present text includes two other projective enumerative computations taken from Salmon’s book; they both fit in the framework of what we would call reduced elimination theory. The first one gives the number of bitangent lines to a surface in $\mathbf{P}^{3}$ through a general point. We generalize it in two ways: to the enumeration of lines in $\mathbf{P}^{n}$ having contact of the form $2p+(n-1)q+\cdots$ with a hypersurface on the one hand, and to the enumeration of bitangent codimension 2 linear spaces on the other hand. While Salmon’s computation and the first generalization do not involve openly the reduced discriminant (thanks to the existence of a cheap substitute in the two indeterminates case), it is essential for the latter generalization. The second computation is that of the degree of the flecnodal curve of a surface $S\subseteq\mathbf{P}^{3}$, which is the locus of points $p$ at which the tangent plane cuts out a curve with one of its two branches at $p$ that has an inflection point. We present here Salmon’s ideas, noting that this computation is carried out in modern standards and in arbitrary dimension in [3].
@@
[…]
ThD thanks Ragni Piene and Israel Vainsencher for their interest, and for suggesting he talks with LB; he also thanks François Lê for sharing some insights on Salmon’s work. LB is grateful to Alexandru Dimca for useful discussions on Milnor number and for indicating the reference [11].
@@
(1.1.1) $\bar{I}$ =$=\bigl{\{}F\in\mathbf{k}[\mathbf{X}]:\forall i=0,\ldots,n,\ \exists N_{i}\ \text{such that}\ X_{i}^{N_{i}}F\in I\bigr{\}}$
@@
In other words, a subscheme $V\subseteq\mathbf{P}^{n}_{\mathbf{k}}$ is defined (scheme-theoretically) by a homogeneous ideal $I\subseteq\mathbf{k}[\mathbf{X}]$ if and only if the saturation $\bar{I}$ equals the homogeneous ideal $I_{V}$ of $V$.
In particular, if $\mathbf{k}$ is a field, a subscheme $V\subseteq\mathbf{P}^{n}_{\mathbf{k}}$ defined by a homogeneous ideal $I$ is empty if and only if the degree 00 piece $\bar{I}_{0}$ is non-zero: this is the homogeneous Nullstellensatz! (see also [6]). On the other hand, the non-emptiness of $V$ is equivalent to the existence of a point in $V(\mathbf{k}^{\prime})$ for some finite field extension $\mathbf{k}^{\prime}$ of $\mathbf{k}$.
When $\mathbf{k}$ is an arbitrary commutative integral domain,222this arguably unconventional notation is used ubiquitously by Jouanolou. the vanishing of $\bar{I}_{0}$ is equivalent to the scheme-theoretic image of the map $V\to\mathrm{Spec}(\mathbf{k})$ being equal to the whole $\mathrm{Spec}(\mathbf{k})$. Indeed, the subscheme of $\mathrm{Spec}(\mathbf{k})$ defined by $\bar{I}_{0}$ coincides as a set with the image of $V\to\mathrm{Spec}(\mathbf{k})$ — this is the proof that projective morphisms are closed —, and moreover $\bar{I}_{0}$ defines the scheme-theoretic image of $V\to\mathrm{Spec}(\mathbf{k})$ (see, e.g., [15, §1] for more details).
@@
Let $\mathbf{k}$ be a commutative ring, and let $u_{j,{\boldsymbol{\alpha}}}\in\mathbf{k}$ for for all $j=1,\ldots,r$ and $|{\boldsymbol{\alpha}}|=d_{j}$. For all $T\in A_{\mathbf{Z}}{}[\mathbf{X}]$, we denote by $T(u_{j,{\boldsymbol{\alpha}}})\in\mathbf{k}[\mathbf{X}]$ the polynomial obtained by evaluating the indeterminates $U_{j,{\boldsymbol{\alpha}}}$ at $u_{j,{\boldsymbol{\alpha}}}\in\mathbf{k}$ for all $j=1,\ldots,r$ and $|{\boldsymbol{\alpha}}|=d_{j}$ in the polynomial $T$.
@@
In other words, the inertia forms for $F_{1},\ldots,F_{r}$ are the homogeneous elements of the saturation of the ideal $(F_{1},\ldots,F_{r})$ in $A_{\mathbf{Z}}{}[\mathbf{X}]$. We denote by $\mathfrak{I}_{\mathbf{Z}}$ the ideal of all degree 00 inertia forms for $F_{1},\ldots,F_{r}$; thus, in the notation introduced in (1.1.1),
@@
Now, in view of the above considerations, we have the following.
@@
This tells us that a given specialization to a field of the polynomials $F_{j}$ defines a non-empty subscheme if and only if all the constants in the saturation of $(F_{1},\ldots,F_{r})$ vanish in this specialization (see also [6]). Note that for all $T\in\mathfrak{I}_{\mathbf{Z}}$, one has $T(u_{j,{\boldsymbol{\alpha}}})\in\mathbf{k}$, since $T$ has degree 00 in $\mathbf{X}$.
@@
In the notation of Paragraph (1.2), let $k$ be an integer in $[\kern-1.72218pt[0,n]\kern-1.72218pt]$, and consider the grading on $A_{\mathbf{Z}}=\mathbf{Z}[U_{j,{\boldsymbol{\alpha}}}]_{j=1,\ldots,n+1,\,|{\boldsymbol{\alpha}}|=d_{j}}$ defined by
@@
The Bezout Theorem, which counts the number of roots of a finite complete intersection scheme in a projective space, and thus is the mother of all statements in projective enumerative geometry, can be deduced from this property. This goes as follows.
@@
(1.10.1) $\mathrm{weight}(U_{j,{\boldsymbol{\alpha}}})=d_{j}-\alpha_{k}.$
@@
in $\mathbf{k}[X_{1},\ldots,X_{n}][T]$. We are abusing notation here, as one should consider instead the two polynomials $F(T,S\,X_{1},\ldots,S\,X_{n})$ and $G(T,S\,X_{1},\ldots,S\,X_{n})$ that are homogeneous in the couple of indeterminates $(S,T)$. Moreover, beware that (1.1.1) above is written “in reverse order” with respect to what we did when discussing the Bezout Theorem, see (1.9.1).
@@
The purpose of reduced elimination theory is to study inertia forms of the truncations $H_{1},\ldots,H_{n+1}$ defined above. Note that, for all $j=1,\ldots,n+1$, the truncation $H_{j}$ is the generic homogeneous polynomial of degree $d_{j}$ with a multiplicity $s_{j}$ zero in the origin $(1:0:...:0)$. The wish for the reduced resultant is that it is a polynomial in the coefficients of $H_{1},\ldots,H_{n+1}$, which vanishes if and only if $H_{1},\ldots,H_{n+1}$ have a non-trivial common root in addition to that of multiplicity $s_{1}\cdots s_{n+1}$ at the origin.
@@
If there is only one integer $j\in[\kern-1.72218pt[1,n+1]\kern-1.72218pt]$ such that $d_{j}>s_{j}$, then the reduced resultant is equal to the resultant of the polynomials $H_{1},\ldots,H_{j-1},\,H_{j+1},\ldots,H_{n+1}$.
@@
Its vanishing on an algebraically closed field $\mathbf{k}$ characterizes those collections of hypersurfaces of $\mathbf{P}^{n}_{\mathbf{k}}$ defined by $h_{1},\ldots,h_{n+1}$ that have a further intersection point, infinitely near or not, besides the origin $(1:0:\ldots:0)$, i.e., those collections such that one of the two following conditions holds:
@@
    the hypersurfaces defined by $h_{1},\ldots,h_{n+1}$ intersect at a point which is different from $(1:0:\ldots:0)$;
@@
    the polynomials $f_{j,s_{j}}$, $j=1,\ldots,n+1$, have a common root in $\mathbf{P}^{n-1}_{\mathbf{k}}$, which means that the tangent cones of the hypersurfaces $\mathrm{V}(h_{1}),\ldots,\mathrm{V}(h_{n+1})$ at the point $(1:0:\ldots:0)$ have a line in common.
@@
and weight 00 to the constants. We find it helpful to visualize this definition as follows:
(1.1.1) $F_{j}=\underbrace{X_{0}^{d_{j}}F_{j,0}}_{\text{coeffs have }\mathrm{weight}\ s_{j}}+\,\cdots\,+\quad\underbrace{X_{0}^{d_{j}-s_{j}+1}F_{j,s_{j}-1}}_{\text{coeffs have }\mathrm{weight}\ 1}\quad+\quad\underbrace{X_{0}^{d_{j}-s_{j}}F_{j,s_{j}}+\cdots+F_{j,d_{j}}}_{\text{coeffs have }\mathrm{weight}\ 0}.$
Note in particular that the indeterminates $U_{j,{\boldsymbol{\alpha}}}$ whose weight equals 00 in this grading are exactly the coefficients of the truncation $H_{j}$ of the polynomial $F_{j}$. The grading (1.8.1) introduced in Proposition (1.8) is a particular case of a Zariski grading (corresponding to $s_{j}=d_{j}$ for all $j$), which explains the terminology “generalized weight properties”.
@@
Observe that, in the Zariski grading: (i) the coefficients of $H_{1},\ldots,H_{n+1}$ all have weight 00, hence $\operatorname{redRes}(H_{1},\ldots,H_{n+1})$ is homogeneous of degree 00; (ii) for all $j=1,\ldots,n+1$, the coefficient of $G_{j}$ in $\mathbf{X}^{\boldsymbol{\alpha}}$ has weight $\alpha_{0}$ hence, by Proposition (1.8), $\mathrm{Res}(G_{1},\ldots,G_{n+1})$ is homogeneous of weight $s_{1}s_{2}\cdots s_{n+1}$. Thus, indeed,
@@
[…]
@@
which gives the result since $d(d-1)$ is always even. ∎
@@
Let $k$ be an integer in $[\kern-1.72218pt[0,n]\kern-1.72218pt]$.
@@
One may use Proposition (2.6.1) to compute the degree of the dual to a smooth hypersurface in $\mathbf{P}^{n+1}$, see Paragraph (3.2).
@@
where each $F_{k}$ is homogeneous of degree $k$ in the indeterminates $X_{1},\ldots,X_{n}$. Given an integer $s\in[\kern-1.72218pt[2,d-1]\kern-1.72218pt]$, we consider the truncation $H$ of $F$ at order $d-s$ with respect to $X_{0}$, defined as in Paragraph (1.2), and set $G$ as in Paragraph (1.2); thus,
@@
The polynomial $H$ has degree $d$, and has valuation $s$ with respect to the indeterminates $X_{1},\ldots,X_{n}$; moreover its partial derivatives with respect to the indeterminates $X_{1},\ldots,X_{n}$ all have degree $d-1$, and valuation $s-1$, so that the reduced resultant
@@
Both discriminants $\mathrm{Disc}(F_{s})$ and $\mathrm{Disc}(F_{d})$ are irreducible as elements of $A_{\mathbf{Z}}{}$. In addition, their vanishing implies the vanishing of the reduced resultant by (1.4). Indeed, the vanishing of $\mathrm{Disc}(F_{d})$ implies the existence of a common root at infinity ($x_{0}=0$) of the polynomial system $\partial_{1}H=\cdots=\partial_{n}H=H=0$. In the same way, the vanishing of $\mathrm{Disc}(F_{s})$ implies the existence of a common root of the polynomial system $\partial_{1}H=\cdots=\partial_{n}H=H=0$ infinitely near to the point $(1:0:\ldots:0)$. We thus conclude that $\mathrm{Disc}(F_{s})$ and $\mathrm{Disc}(F_{d})$, which are coprime as they are irreducible and do not depend on the same coefficients of $F$, both divide $\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)$, which ends the proof. ∎
@@
specializing the variable $X_{0}$ to 00 in (2.2.1), we get that
@@
The reduced discriminant of $F$ with respect to the truncation at order $d-s$ for the indeterminate $X_{0}$, denoted by $\operatorname{redDisc}^{s}_{d}(H)$, or simply $\operatorname{redDisc}(H)$, is defined by the equality
@@
(observe that $\partial_{0}H$ is of degree $d-1$ and valuation $s$, and $\left.H\right|_{X_{0}=0}=F_{d}$). We conclude the proof of (2.4.1) before justifying the two above equalities. The Zariski formula together with the property of resultants under permutation of the polynomials [15, §5.8] yields the equality:
@@
which follows from the Zariski formula. Indeed, let $F_{1},\ldots,F_{n+1}$ be generic polynomials as above, and denote by $H_{i}$ and $G_{i}$ their decompositions corresponding to the truncation at order $d_{i}-s_{i}$ for all $i$. On the one hand, we have (observe that $F_{n+1}$ and $X_{0}F_{n+1}$ have the same “$G$”, namely $G_{n+1}$):

### v3

# Generalized weight properties of resultants and discriminants, and applications to projective enumerative geometry

@@
Generalized weight properties of resultants and discriminants, and applications to projective enumerative geometry
by Laurent Busé and Thomas Dedieu footnotetext: This text is from a series originated in the seminar Degenerations and enumeration of curves on surfaces, held at the University of Roma “Tor Vergata” in the years 2015–2017. See https://www.math.univ-toulouse.fr/˜tdedieu/#enumTV for the full collection.
In his wonderful book [18], first published in 1862, Salmon casually gives the leading term of some Taylor expansion of the discriminant of a plane algebraic curve. Consider the curve $\mathrm{V}(F)\subseteq\mathbf{P}^{2}$; in a suitable system of homogeneous coordinates (X:Y:Z)(X\!:\!Y\!:\!Z), its equation $F$ has the form
where each polynomial $F_{k}(X,Y)$ is homogeneous of degree $k$ in $X,Y$, and $T,A,B,C$ are coefficients of $F$. Salmon states, without any explanation or reference, that the discriminant of $F$ is of the form
@@
where $\Phi$ is “the discriminant when $T$ vanishes” (and $\Psi$ is some polynomial in the coefficients of $F$) [18, § 605]. This note arose as an attempt to understand and prove this formula.
We introduce the reduced discriminant of a hypersurface (or homogeneous polynomial), of which $\Phi$ is an instance, it turns out. It is deeply linked with the reduced resultant, introduced by Zariski in [21], in 1936 long after Salmon. Of course, when $T$ is zero, $\mathrm{V}(F)$ is singular at the point $(0:0:1)$, no matter the other coefficients of $F$, and correspondingly $\mathrm{Disc}(F)$ vanishes identically. The polynomial $\Phi$, in fact, vanishes at those values of the coefficients of $F$ for which the curve $\mathrm{V}(F)$ has some additional singularity besides the point $(0:0:1)$. The reduced discriminant generalizes this to the case when $F$ has a singular point of arbitrary multiplicity $s$. Somehow built in the theory is a general form of Formula (a). We show how to get (a) specifically (in fact with a normalization factor $\frac{1}{2}$, which is probably irrelevant for Salmon), and generalize the formula for hypersurfaces of arbitrary dimension, see Section 2.3. This is certainly what Salmon has in mind, although arguably only for $s=2$, and possibly only empirically. It has since then been apparently completely forgotten; we have only found a faint trace of these ideas in Salmon’s works, see Paragraph (2.12). We point out that the reduced discriminant is a particular instance of the theory of toric discriminants studied in [12], which is more general in terms of the family of hypersurfaces that can be considered. The methodology that we develop here in the specific setting of the reduced discriminant is different from the theory of toric discriminants, and we found it particularly well adapted, if not necessary, to tackle Salmon’s formula.
Salmon then uses Formula (a) to derive various enumerative quantities for surfaces $S\subseteq\mathbf{P}^{3}$. In particular, he computes the number of bitangent planes passing through a fixed general point $p\in\mathbf{P}^{3}$. We explain his method, which involves a remarkable trick in elimination theory, and generalize it for hypersurfaces of arbitrary dimension. Salmon’s strategy is to consider a pencil of planes with center a line tangent to the surface $S$ at some point $p^{\prime}$; this pencil contains a finite number of planes tangent to $S$, among which the tangent plane at $p^{\prime}$ counts with multiplicity $2$ in general, and with greater multiplicity if it has some special feature, e.g., if it is a bitangent plane. In an appropriate setting, this multiplicity is the order in $T$ of the polynomial in (a), and the game is to understand the conditions on the point $p^{\prime}$ that make this multiplicity jump. In the upshot we get, indirectly from the reduced discriminant $\Phi$, a curve in $S$ parametrizing those points $p^{\prime}$ such that the tangent plane at $p^{\prime}$ is a bitangent plane; following Salmon, we call it the node-couple curve. There are other well-known ways to compute the number of bitangent planes to a surface in $\mathbf{P}^{3}$ (see, e.g., [II]). The present method has the advantage of being core level: it is only a study of homogeneous polynomials and their discriminants.
@@
The present text includes two other projective enumerative computations taken from Salmon’s book; they both fit in the framework of what we would call reduced elimination theory. The first one gives the number of bitangent lines to a surface in $\mathbf{P}^{3}$ through a general point. We generalize it in two ways: to the enumeration of lines in $\mathbf{P}^{n}$ having contact of the form $2p+(n-1)q+\cdots$ with a hypersurface on the one hand, and to the enumeration of bitangent codimension 2 linear spaces on the other hand. While Salmon’s computation and the first generalization do not involve openly the reduced discriminant (thanks to the existence of a cheap substitute in the two indeterminates case), it is essential for the latter generalization. The second computation is that of the degree of the flecnodal curve of a surface $S\subseteq\mathbf{P}^{3}$, which is the locus of points $p$ at which the tangent plane cuts out a curve with one of its two branches at $p$ that has an inflection point (equivalently, there is a line having contact order with $S$ at $p$ at least $4$). We present here Salmon’s ideas, noting that this computation is carried out in modern standards and in arbitrary dimension in [3].
@@
We put particular focus on the homogeneity properties of the resultants and discriminants, with respect to various gradings. They are indeed essential for the enumerative applications, and we give numerous illustrations, including a proof of Bezout’s theorem and the computation of the degree of the dual of a smooth hypersurface. Beyond that, they are central in the theory of reduced resultants and discriminants, see Theorem (1.3) and Corollary (2.13), and indeed the key to Formula (a) and its generalizations.
ThD thanks Ragni Piene and Israel Vainsencher for their interest, and for suggesting he talk with LB; he also thanks François Lê for sharing some insights on Salmon’s work. LB is grateful to Alexandru Dimca for useful discussions on Milnor number and for indicating the reference [11]. We both thank the referees for their helpful observations.
@@
(0.2) Definition.
Given a grading $\mathbf{k}[\mathbf{X}]=\bigoplus_{d\geqslant 0}\mathbf{k}[\mathbf{X}]_{d}$, the order of a polynomial $F\in\mathbf{k}[\mathbf{X}]$ is the largest integer $e$ such that $F\in\bigoplus_{d\geqslant e}\mathbf{k}[\mathbf{X}]_{d}$.
We will usually consider gradings obtained by assigning a weight $w_{i}$ to each indeterminate $X_{i}$: the degree $d$ homogeneous piece $\mathbf{k}[\mathbf{X}]_{d}$ is then generated by all monomials $\mathbf{X}^{\boldsymbol{\alpha}}$ such that $\sum_{i=0}^{n}\alpha_{i}w_{i}=d$.
Given a subset $I\subseteq\{1,\ldots,n\}$, the order of a polynomial $F$ in the indeterminates $X_{i}$, $i\in I$, is the order of $F$ with respect to the grading obtained by assigning to $X_{i}$ the weight $1$ if $i\in I$, and the weight 0 if $i\not\in I$.
Beware that $\mathbf{k}$ itself will often be a polynomial ring, and we will consider gradings and orders both on $\mathbf{k}[\mathbf{X}]$ and on $\mathbf{k}$ itself.
@@
(1.1.1) $\bar{I}$ $=\bigl\{F\in\mathbf{k}[\mathbf{X}]:\forall i=0,\ldots,n,\ \exists N_{i}\ \text{such that}\ X_{i}^{N_{i}}F\in I\bigr\}$
@@
In other words, a subscheme $V\subseteq\mathbf{P}^{n}_{\mathbf{k}}$ is defined (scheme-theoretically) by a homogeneous ideal $I\subseteq\mathbf{k}[\mathbf{X}]$ if and only if the saturation $\bar{I}$ equals the homogeneous ideal $I_{V}$ of $V$. For example, for all $F\in\mathbf{k}[\mathbf{X}]$, the two ideals $(F)$ and $(X_{0}F,\ldots,X_{n}F)$ define the same subcheme of $\mathbf{P}^{n}_{\mathbf{k}}$ (and we recommend that the reader check it for themselves).
In particular, if $\mathbf{k}$ is a field, a subscheme $V\subseteq\mathbf{P}^{n}_{\mathbf{k}}$ defined by a homogeneous ideal $I$ is empty if and only if the degree 0 piece $\bar{I}_{0}$ is non-zero: this is the homogeneous Nullstellensatz! (see also [6]). On the other hand, the non-emptiness of $V$ is equivalent to the existence of a point in $V(\mathbf{k}^{\prime})$ for some finite field extension $\mathbf{k}^{\prime}$ of $\mathbf{k}$.
When $\mathbf{k}$ is an arbitrary commutative integral domain,111this arguably unconventional notation is used ubiquitously by Jouanolou. the vanishing of $\bar{I}_{0}$ is equivalent to the scheme-theoretic image of the map $V\to\mathrm{Spec}(\mathbf{k})$ being equal to the whole $\mathrm{Spec}(\mathbf{k})$. Indeed, the subscheme of $\mathrm{Spec}(\mathbf{k})$ defined by $\bar{I}_{0}$ coincides as a set with the image of $V\to\mathrm{Spec}(\mathbf{k})$ — this is the proof that projective morphisms are closed —, and moreover $\bar{I}_{0}$ defines the scheme-theoretic image of $V\to\mathrm{Spec}(\mathbf{k})$ (see, e.g., [15, §1] for more details).
@@
Let $\mathbf{k}$ be a commutative ring, and let $u_{j,{\boldsymbol{\alpha}}}\in\mathbf{k}$ for all $j=1,\ldots,r$ and $|{\boldsymbol{\alpha}}|=d_{j}$. For all $T\in A_{\mathbf{Z}}{}[\mathbf{X}]$, we denote by $T(u_{j,{\boldsymbol{\alpha}}})\in\mathbf{k}[\mathbf{X}]$ the polynomial obtained by evaluating the indeterminates $U_{j,{\boldsymbol{\alpha}}}$ at $u_{j,{\boldsymbol{\alpha}}}\in\mathbf{k}$ for all $j=1,\ldots,r$ and $|{\boldsymbol{\alpha}}|=d_{j}$ in the polynomial $T$.
@@
In other words, the inertia forms for $F_{1},\ldots,F_{r}$ are the elements of the saturation of the ideal $(F_{1},\ldots,F_{r})$ in $A_{\mathbf{Z}}{}[\mathbf{X}]$. We denote by $\mathfrak{I}_{\mathbf{Z}}$ the ideal of all degree 0 inertia forms for $F_{1},\ldots,F_{r}$; thus, in the notation introduced in (1.1.1),
@@
Note that this is a homogeneous ideal with respect to the standard grading on $A_{\mathbf{Z}}$. In view of the above considerations, we have the following.
@@
This tells us that a given specialization to a field of the polynomials $F_{j}$ defines a non-empty subscheme if and only if all the constants in the saturation of $(F_{1},\ldots,F_{r})$ vanish in this specialization (see also [6]). Note that for all $T\in\mathfrak{I}_{\mathbf{Z}}$, one has $T(u_{j,{\boldsymbol{\alpha}}})\in\mathbf{k}$, since $T$ has degree 0 in $\mathbf{X}$.
(1.5) Example.
Set $n=1$, and let $F_{1}=U_{1,0}X_{0}+U_{1,1}X_{1}$ and $F_{2}=U_{2,0}X_{0}+U_{2,1}X_{1}$. Then $R=U_{1,0}U_{2,1}-U_{1,1}U_{2,0}$ is a degree 0 inertia form of $F_{1}$ and $F_{2}$, as indeed
$X_{0}R=U_{2,1}F_{1}-U_{1,1}F_{2}\quad\text{and}\quad X_{1}R=U_{1,0}F_{2}-U_{2,0}F_{1}.$
We thus recover the usual determinant of two linear forms. We recommend that the reader come back to this example when we introduce various gradings on $A_{\mathbf{Z}}$ later on.
@@
In the notation of Paragraph (1.2), let $k$ be an integer in $\kern-1.72218pt[0,n]\kern-1.72218pt]$, and consider the grading on $A_{\mathbf{Z}}=\mathbf{Z}[U_{j,{\boldsymbol{\alpha}}}]_{j=1,\ldots,n+1,\,|{\boldsymbol{\alpha}}|=d_{j}}$ defined by
@@
Bezout Theorem, which counts the number of roots of a finite complete intersection scheme in a projective space, and thus is the mother of all statements in projective enumerative geometry, can be deduced from this property. This goes as follows.
@@
($1.1$) $\mathrm{weight}(U_{j,{\boldsymbol{\alpha}}})=d_{j}-\alpha_{k}.$
@@
in $\mathbf{k}[X_{1},\ldots,X_{n}][T]$. We are abusing notation here, as one should consider instead the two polynomials $F(T,S\,X_{1},\ldots,S\,X_{n})$ and $G(T,S\,X_{1},\ldots,S\,X_{n})$ that are homogeneous in the couple of indeterminates $(S,T)$. Moreover, beware that (1.1.1) above is written “in reverse order” with respect to what we did when discussing Bezout Theorem, see (1.10.1).
@@
The purpose of reduced elimination theory is to study inertia forms of the truncations $H_{1},\ldots,H_{n+1}$ defined above. Note that, for all $j=1,\ldots,n+1$, the truncation $H_{j}$ is the generic homogeneous polynomial of degree $d_{j}$ with a multiplicity $s_{j}$ zero in the origin (1:0::0)(1:0:...:0). The wish for the reduced resultant is that it is a polynomial in the coefficients of $H_{1},\ldots,H_{n+1}$, which vanishes if and only if $H_{1},\ldots,H_{n+1}$ have a non-trivial common root in addition to that of multiplicity $s_{1}\cdots s_{n+1}$ at the origin.
@@
If there is only one integer $j\in[\kern-1.72218pt[1,n+1]\kern-1.72218pt]$ such that $d_{j}>s_{j}$, then the reduced resultant is equal to the resultant of the polynomials $H_{1},\ldots,H_{j-1},\,H_{j+1},\ldots,H_{n+1}$, as polynomials in $X_{1},\ldots,X_{n}$ only.222if $s_{j}=d_{j}$, then $H_{j}=F_{j,d_{j}}$ is a polynomial in $X_{1},\ldots,X_{n}$ only; geometrically, the polynomial $H_{j}$ defines a cone in $\mathbf{P}^{n}$ over the hypersurface defined by $F_{j,d_{j}}$ in $\mathbf{P}^{n-1}$.
@@
Its vanishing on an algebraically closed field $\mathbf{k}$ characterizes those collections of hypersurfaces of $\mathbf{P}^{n}_{\mathbf{k}}$ defined by $h_{1},\ldots,h_{n+1}$ that have a further intersection point, infinitely near or not, besides the origin $p_{0}=(1\!:\!0\!:\!\ldots\!:\!0)$, i.e., those collections such that one of the two following conditions holds:
@@
    the hypersurfaces defined by $h_{1},\ldots,h_{n+1}$ intersect at a point which is different from $p_{0}$;
@@
    the polynomials $f_{j,s_{j}}$, $j=1,\ldots,n+1$, have a common root in $\mathbf{P}^{n-1}_{\mathbf{k}}$, which means that the tangent cones of the hypersurfaces $\mathrm{V}(h_{1}),\ldots,\mathrm{V}(h_{n+1})$ at $p_{0}$ have a line in common; equivalently, the proper transforms of $\mathrm{V}(h_{1}),\ldots,\mathrm{V}(h_{n+1})$ in the blow-up of $\mathbf{P}^{n}_{\mathbf{k}}$ at $p_{0}$ have a point in common which lies on the exceptional divisor.
@@
and weight 0 to the constants. We find it helpful to visualize this definition as follows:
[…]
Note in particular that the indeterminates $U_{j,{\boldsymbol{\alpha}}}$ whose weight equals 0 in this grading are exactly the coefficients of the truncation $H_{j}$ of the polynomial $F_{j}$. The grading (1.9.1) introduced in Proposition (1.9) is a particular case of a Zariski grading (corresponding to $s_{j}=d_{j}$ for all $j$), which explains the terminology “generalized weight properties”.
@@
Observe that, in the Zariski grading: (i) the coefficients of $H_{1},\ldots,H_{n+1}$ all have weight 0, hence $\operatorname{redRes}(H_{1},\ldots,H_{n+1})$ is homogeneous of degree 0; (ii) for all $j=1,\ldots,n+1$, the coefficient of $G_{j}$ in $\mathbf{X}^{\boldsymbol{\alpha}}$ has weight $\alpha_{0}$ hence, by Proposition (1.9), $\mathrm{Res}(G_{1},\ldots,G_{n+1})$ is homogeneous of weight $s_{1}s_{2}\cdots s_{n+1}$. Thus, indeed,
@@
In this Section we give a rigorous proof of formula (a). This is done by introducing the concept of reduced discriminant. We begin with a quick recap on the ordinary discriminant of a hypersurface, following [5, §4]; see also [8] and [12, Chapter 13, §D].
@@
which gives the result since d(d1)d(d-1) is always even. ∎
@@
Let $k$ be an integer in $\kern-1.72218pt[0,n]\kern-1.72218pt]$.
@@
One may use (2.6.1) to compute the degree of the dual to a smooth hypersurface in $\mathbf{P}^{n+1}$, see Paragraph (3.2).
@@
where each $F_{k}$ is homogeneous of degree $k$ in the indeterminates $X_{1},\ldots,X_{n}$. Given an integer $s\in[\kern-1.72218pt[2,d-1]\kern-1.72218pt]$, we consider the truncation $H$ of $F$ at order dsd-s with respect to $X_{0}$, defined as in Paragraph (1.2), and set $G$ as in Paragraph (1.2); thus,
@@
The polynomial $H$ is homogeneous of degree $d$, and has order $s$ in $X_{1},\ldots,X_{n}$ (recall Definition (0.2)). Moreover, its partial derivatives with respect to the indeterminates $X_{1},\ldots,X_{n}$ all have degree d1d-1, and order s1s-1 in $X_{1},\ldots,X_{n}$. Therefore, the reduced resultant
@@
Both discriminants $\mathrm{Disc}(F_{s})$ and $\mathrm{Disc}(F_{d})$ are irreducible as elements of $A_{\mathbf{Z}}{}$, and therefore coprime because they do not depend on the same coefficients of $F$. To prove the proposition, it thus suffices to prove that the vanishing of $\mathrm{Disc}(F_{s})$ or $\mathrm{Disc}(F_{d})$ implies the vanishing of the reduced resultant. This follows from (1.4), as we shall now explain.
The vanishing of $\mathrm{Disc}(F_{d})$ at $f$ (a specialization of $F$ with coefficients in an algebraically closed field $\mathbf{k}$) implies the existence of a common root at infinity ($x_{0}=0$) of the polynomials $h,\partial_{1}h,\ldots,\partial_{n}h$ (with $h$ the specialization of $H$ corresponding to $f$). Indeed, if $\mathrm{Disc}(F_{d})$ vanishes at $f$, then by Theorem (2.4) there exists a non-trivial zero $\vphantom{x}{}^{\flat}\kern-1.00006ptx\in\mathbf{k}^{n}$ of $f_{d},\partial_{1}f_{d},\ldots\partial_{n}f_{d}$, and then $(0,\vphantom{x}^{\flat}\kern-1.00006ptx)\in\mathbf{k}^{n+1}$ is a non-trivial zero of $h,\partial_{1}h,\ldots,\partial_{n}h$, hence the reduced resultant vanishes at $f$ by (a) of (1.4).
Similarly, the vanishing of $\mathrm{Disc}(F_{s})$ at $f$ implies the existence of a common non-trivial zero of $h,\partial_{1}h,\ldots,\partial_{n}h$ infinitely near the point $p_{0}=(1\!:\!0\!:\!\ldots\!:\!0)$: indeed, $f_{s},\partial_{1}f_{s},\ldots,\partial_{n}f_{s}$ define the respective tangent cones at $p_{0}$ of the hypersurfaces $V(h),V(\partial_{1}h),\ldots,V(\partial_{n}h)$, and the vanishing of $\mathrm{Disc}(F_{s})$ at $f$ implies the existence of a non-trivial zero $\vphantom{x}{}^{\flat}\kern-1.00006ptx\in\mathbf{k}^{n}$ of $f_{s},\partial_{1}f_{s},\ldots\partial_{n}f_{s}$. Thus, by (b) of (1.4), the reduced resultant vanishes at $f$ if $\mathrm{Disc}(F_{s})$ vanishes at $f$. ∎
@@
specializing the variable $X_{0}$ to 0 in (2.2.1), we get that
@@
The reduced discriminant of $F$ with respect to the truncation at order dsd-s for the indeterminate $X_{0}$, denoted by $\operatorname{redDisc}^{s}_{d}(H)$, or simply $\operatorname{redDisc}(H)$, is defined by the equality
@@
(observe that $\partial_{0}H$ has degree d1d-1 and order $s$ in $X_{1},\ldots,X_{n}$, and $\left.H\right|_{X_{0}=0}=F_{d}$). We conclude the proof of (2.4.1) before justifying the two above equalities. The Zariski formula together with the property of resultants under permutation of the polynomials [15, §5.8] yields the equality:
@@
which follows from the Zariski formula. Indeed, let $F_{1},\ldots,F_{n+1}$ be generic polynomials as above, and denote by $H_{i}$ and $G_{i}$ their decompositions corresponding to the truncation at order $d_{i}-s_{i}$ for all $i$. On the one hand, we have (observe that $F_{n+1}$ and $X_{0}F_{n+1}$ have the same “GG”, namely $G_{n+1}$):